# Overcoming the Zero-Rate Hashing Bound with Holographic Quantum Error Correction

### v2

@@ Abstract
A crucial insight for practical quantum error correction is that different types of errors, such as single-qubit Pauli operators, typically occur with different probabilities. Finding an optimal quantum code under such biased noise is a challenging problem, related to finding the (generally unknown) maximum capacity of the corresponding noisy channel. A benchmark for this capacity is given by the hashing bound, describing the performance of random stabilizer codes, which leads to the challenge of finding codes that reach or exceed this bound while also being efficiently decodable. In this work, we show that asymptotically zero-rate holographic codes, built from hyperbolic tensor networks that model holographic bulk/boundary dualities, fulfill both conditions. Of the five holographic code models considered, all are found to reach the hashing bound in some bias regime and one, the holographic surface-code fragment, appears to even exceed the capacity of previously known codes in the 2-Pauli-dominated noise regime. In addition, we consider Clifford deformations that allow all considered codes to reach the hashing bound for 1-Pauli-dominated noise as well. Our results thus establish that holographic codes, which were previously shown to possess efficient tensor-network decoders, also exhibit competitive thresholds under biased noise.
**footnotetext: These authors contributed equally to this work.
@@ I Introduction
Understanding and mitigating the impact of quantum noise is crucial for quantum computation and communication. The most common way to characterize quantum noise is through quantum channels, which provide rigorous descriptions for how a quantum state is transformed or transmitted in the presence of errors. The maximum number of qubits that can be transmitted reliably in a noisy environment is known as the quantum channel capacity (QCC) [1, 2, 3, 4, 5, 6], which provides an important theoretical limit for quantum computation and communication tasks. However, determining the precise value of the QCC has been incredibly challenging except for a few isolated instances [7]. In particular, the QCC is not accurately known even for a channel as ubiquitous as the depolarizing channel. A large part of the difficulty is related to the potential superadditivity of coherent information, which hinders a direct efficient evaluation of the quantum capacity when the channel acts independently and identically over many qubits. As a result, finding lower and upper bounds of the quantum capacity constitute important progress in quantum Shannon theory [8, 1].
Aside from theoretical limits, it is also of practical interest to design classes of quantum error correction (QEC) codes that can come arbitrarily close to the QCC, thus maximizing the protection of information during noisy transmission or computation. The leading method for generating such practical QEC codes is to consider stabilizer codes, which are closely connected to classical linear codes [9, 10, 11]. However, unlike random classical linear codes which are typically capacity-achieving for symmetric channels, random quantum stabilizer codes do not generally achieve the optimal transmission rate for their analogous quantum counterparts such as the depolarizing channel or other Pauli channels. Instead, they attain a theoretical lower bound on QCC called the hashing bound [12, 13, 8]. Currently, there remains a large gap between the theoretical QCC upper bound and the hashing lower bound. In fact, only tiny improvements over the hashing bound have been made over the years for Pauli channels [2, 14, 15]. Furthermore, optimal decoding of random stabilizer codes is #$\#P$-complete [16], severely limiting their applicability. Therefore, it is of both theoretical and practical interest in quantum coding theory to identify efficiently decodable codes that can attain or even exceed the hashing bound. A class of such codes will not only provide new approaches for obtaining tighter QCC lower bounds but also inspire proposals for more efficient strategies of error correction.
Along this direction, a significant amount of work has been carried out over the past few decades, which has led to a number of QEC codes that approach the hashing bound, e.g. quantum turbo codes and related variations [17, 18, 19]. More recently, certain topological codes (e.g. Clifford-deformed topological codes such as the XZZX surface code) and certain concatenated codes built from various-sized repetition codes have also been able to attain or exceed the error threshold set by the hashing limit [20, 21, 22, 23, 24, 25, 26, 27].
In this work, we demonstrate that asymptotically zero-rate holographic quantum codes, a relatively recent code class derived from properties of the anti-de Sitter / conformal field theory correspondence (AdS/CFT) [28, 29, 30, 31, 32, 33, 34, 35], can achieve and supersede the hashing bound for various Pauli channels of interest. Using the tensor-network decoder described in [36, 37], we study zero-rate versions of five different holographic codes: the Harlow-Preskill-Pastawski-Yoshida (HaPPY) code [33]; the tailored $\llbracket 7,1,3\rrbracket$ code [38]; the holographic Steane code [39]; the holographic $\llbracket 6,1,3\rrbracket$ code [36]; and the holographic surface-code fragment (SCF) [40].
@@
Our main results can be stated as follows. In the limit of pure 2-Pauli noise channels, the thresholds we have calculated for zero-rate SCF and $\llbracket 6,1,3\rrbracket$ holographic codes exceed current state-of-the-art results [27, 15, 2]. To our knowledge, no other class of quantum code has thus far been able to supersede the hashing limit for pure 2-Pauli noise. In fact, all of the holographic codes tested in this limit either come within $2\%$ of, attain, or overcome the hashing bound completely. Moreover, all of the codes tested were found to either attain or overtake the hashing bound for numerous distinct Pauli channels. Finally, we also apply the method of Clifford deformations from [21, 20] to the holographic Steane code, demonstrating that the threshold of the code can be drastically modified for depolarizing, pure X, and pure Z noise channels. Our results provide strong evidence that holographic codes possess very high resilience against various types of Pauli noise channels and can be tuned accordingly. In sharp contrast with topological codes, maximum likelihood decoding of holographic codes and the computation of weight enumerator polynomials are possible in polynomial time, even with exact tensor network contractions [36, 41], further bolstering their practical potential for both quantum error correction and for numerical studies of QCC bounds.
@@
The remainder of this article is organized as follows: we review in the background (Section II) the basics of holographic codes and their properties for practical QEC in Section II.1, as well as an explanation for how tensor-network decoding functions (Section II.2). In Section III, we begin by discussing first the experimental setup for our simulations, the codes tested, and the biased-noise points considered (Section III.1); in Section III.2, we present all of the results of our study. Finally, we close with a discussion of future directions and implications of our work (Section IV), followed by a discussion of possible practical applications for holographic codes (Section IV.1).
@@ II.1 Holographic Quantum Error Correction
Figure 1: The holographic tensor-network codes considered in this work: (a) The Harlow-Preskill-Pastawski-Yoshida (HaPPY) code [33] whose pentagon-hexagon geometry also underlies the holographic surface-code fragment (SCF) model [40], (b) the holographic $\llbracket 6,1,3\rrbracket$ code [36], and (c) the holographic Steane code [39]. All are defined as tensor network contractions of copies of a fixed $q$-leg tensor on a hyperbolic tiling, with the central tensor used as a seed tensor for the encoding isometry of a $\llbracket q-1,1,d\rrbracket$ code with some distance $d$. The remaining tensors have all $q$ legs contracted in the plane, leading to a larger encoding map between one logical qubit (central red leg) and the boundary physical qubits (open black legs). Here we depict the codes with $L=2$ layers of edge inflation. In the $L\to\infty$ limit, the rate of each code goes to zero.
Holographic quantum codes at their core are subsystem stabilizer codes [42, 43, 44]. Stabilizer codes themselves are defined as codes for which the logical operators $\mathsf{L}$ and codeword stabilizers $\mathsf{S}$ consist of elements $P_{i}\in\mathsf{P}^{n}$, where $\mathsf{P}^{n}$ is the $n$-qubit Pauli group, $P_{i}$ takes the form $P_{i_{1}}\otimes\cdots\otimes P_{i_{n}}$, and $P\in\{I,X,Y,Z\}$, i.e., the single-qubit Pauli operators. In subsystem stabilizer codes such as holographic codes, one typically can choose to subdivide the physical Hilbert space as $\mathsf{H}_{p}=\mathsf{H}\otimes\mathsf{\bar{H}}$, where $\mathsf{H}$ represents the code subspace. It is also given that $\mathsf{H}$ can be decomposed as $\mathsf{H}=\mathsf{H}_{L}\otimes\mathsf{H}_{G}$, where $\mathsf{H}_{L}$ and $\mathsf{H}_{G}$ represent the logical and gauge subsystems of the code subspace, respectively. Holographic codes were first treated as such subsystem stabilizer codes in [45, 43, 46]. It is also wholly possible to consider finite-rate versions of holographic codes, wherein individual seed tensors can be specified with logical-qubit implantation. In this work, we consider only zero-rate versions of holographic codes, wherein only the central logical index remains (Figure 1). Additionally, we mention briefly that it is possible to construct non-Abelian stabilizer, non-stabilizer, or even approximate quantum codes in the holographic context [47, 48, 45, 49, 50]; however, we shall limit ourselves to stabilizer versions of holographic codes in the present work.
@@
Originally developed to model the AdS/CFT correspondence [28, 29, 30, 51, 52], holographic quantum error correction involves encoding maps from bulk (logical) degrees of freedom on a $d{+}1$-dimensional hyperbolic space to $d$-dimensional boundary (physical) degrees of freedom. While in full AdS/CFT the bulk and boundary degrees of freedom are associated with weakly-coupled quantum gravity on an asymptotically anti-de Sitter (AdS) background and strongly-coupled conformal field theory (CFT), respectively, aspects of this duality for $d=1$ can be captured by simple tensor network codes [31, 33, 34, 35].
@@
From a practical perspective, there are many reasons to find the concept of a holographic quantum code appealing. Firstly, holographic codes are relatively simple to construct and scale from small, atomized examples to larger example codes. As a direct consequence of the quantum LEGO formalism [53], concepts such as code structure and transversal logical operations are straightforward to intuit. Secondly, the boundary of holographic codes exhibits a quasiperiodic self-similarity [54, 55], allowing one to rescale the number of physical and logical qubits using local inflation rules that also determine the code’s rate and distance scaling [37, 56], both of which fare better than in constructions of topological quantum codes [57, 58, 59], as well as in many of the recent proposals for quantum low-density parity-check (qLDPC) codes [60]. Thirdly, it is known that most seed tensors for holographic codes can be easily mapped to graph states, implying that efficient preparation schemes likely exist [61, 62, 63, 64]. Lastly, and perhaps most consequentially, it has been shown that holographic codes exhibit high resilience against various noise channels. Indeed, it was first shown in [33] that holographic quantum error correction codes exhibit high thresholds against the quantum erasure channel; subsequently, several works have shown the capability of certain codes constructions to protect against depolarizing and 1-Pauli noise, potentially as well as topological codes [37, 39, 40, 36, 65, 56]. To our knowledge, a systematic biased-noise threshold study for even zero-rate holographic codes has not been performed, let alone for their constant-rate versions111 [41] discusses biased noise in finite-rate HaPPY pentagon codes, but it only captures the non-detectable error probability for a code of fixed length.. As such, our work takes the first step towards understanding holographic codes under more generalized noise channels.
@@
The main ingredient of a holographic quantum code is the seed tensor defining the encoding map for a single tile. Here we consider the zero-rate case with a logical qubit only on the central tile; on the remaining tiles all tensor legs are planar, effectively dispersing the logical information towards the tiling boundary. Note that in generic holographic codes, each tile may hold a logical qubit, leading to a nonzero asymptotic rate as the number $L$ of layers is increased. It was recently proven that holographic codes are an instance of generalized code concatenation [66, 67], and, while copies of the same seed tensor are typically used, one may also employ several different seed tensors in similar spirit to heterogeneous concatenation methods for tree-style concatenated codes [68, 69, 67].
II.2 Tensor-Network Decoding

### v3

# Biased-Noise Thresholds of Zero-Rate Holographic Codes with Tensor-Network Decoding

@@ Abstract
A crucial insight for practical quantum error correction is that different types of errors, such as single-qubit Pauli operators, typically occur with different probabilities. Finding an optimal quantum code under such biased noise is a challenging problem, related to the (generally unknown) maximum capacity of the corresponding noisy channel. A benchmark for this capacity is given by the hashing bound, which describes the performance of random stabilizer codes and leads to the matter of identifying codes that come close to the bound while also being efficiently decodable. In this work, we perform the first comprehensive analysis of asymptotically zero-rate holographic codes under biased noise. We show that many representatives from such models of this code class fulfill both the channel optimality and efficient decoding guarantees for tensor-network codes. In fact, all holographic codes tested were found to reach the hashing bound in some bias regime, while several built from the $\llbracket 5,1,2\rrbracket$ surface code and $\llbracket 6,1,3\rrbracket$ code exceed state-of-the-art code performance in the 2-Pauli noise regime. Furthermore, we consider Clifford deformations which allow all considered codes to reach the hashing bound for 1-Pauli noise as well. Our results establish that holographic codes, which were previously shown to possess efficient tensor-network decoders, also exhibit competitive thresholds under biased noise.
22footnotetext: These authors contributed equally to this work.33footnotetext: Corresponding author: matt.steinberg3@gmail.com
@@ I Introduction
Holographic codes are quantum codes conventionally studied as toy models of the Anti-de Sitter/Conformal Field Theory (AdS/CFT) correspondence [witten, maldacena, klebanov, almheiri2015bulk, harlow2017ryutakayanagi, happy_paper, Harlow:2018fse, holographic_codes_topical_review] where they have led to remarkable insights in quantum gravity. However, because their relevance in quantum gravity is mostly theoretical [uberholography, harlow2017ryutakayanagi], very little is known about their practical error correction capabilities in contrast to leading benchmarks of quantum error correction, such as the surface code. In addition, recent advances in tensor network codes [farrelly_tn_codes] and the quantum LEGO formalism [cao_quantum_lego] have brought them closer to a practically viable regime. For example, leveraging their flexibility in the modular construction, holographic codes with high encoding rate and greatly exceeding the $O(n^{1/2})$ distance scaling have been proposed [evenbly_codes2] alongside schemes for implementing universal fault-tolerant logic based on heterogeneous code constructions [heterogeneous_holo_qrm]. The hyperbolic geometry of holographic codes also implies that such codes admit efficient optimal tensor network decoders for all i.i.d. single qubit errors [cao2024quantum] since their tensor networks can be contracted in polynomial time even without bond truncation. Under such decoders, they have been shown to yield a code capacity threshold that are as high as that of the surface code [farrelly_tn_codes] under depolarizing noise where all Pauli errors occur with equal probability.
Although there remain many open questions before holographic codes can be fairly assessed at a practical level, e.g. determining the fault-tolerant threshold, there are properties that are poorly understood even at the code capacity level, where syndrome extractions are perfect. One such open question is the performance of the code when the error channels are biased. Biased noise is found in many physical quantum devices wherein certain Pauli errors such as phase flips, occur more frequently than bit flips. The presence of bias can also lead to better error correction with superior threshold and effective distances [sahay2023high, xzzx_surface_code, tailored_513_xzzx_codes, clifford_deformed_surface]. In certain topological codes (e.g. Clifford-deformed topological codes such as the XZZX surface code) the threshold under biased noise has also been shown to attain the error threshold set by the hashing limit, which sets a target benchmark for good quantum error correcting codes that can be used for efficient information transmission under a noisy channel. In holographic codes, however, there is very little understanding of its performance against biased error [cao2024quantum, parallel_decoding_tn_codes].
In this work, we provide the first characterization of holographic codes under biased noise. We study common classes of holographic codes and their deformations to demonstrate that asymptotically zero-rate holographic quantum codes admit high code capacity thresholds comparable to leading benchmarks under biased noise and can either achieve or supersede the hashing bound for various Pauli channels of interest. The possibility of holographic codes as an alternative class of code constructions to exceed the hashing bound is also of independent interest, as a central problem in quantum coding theory is to identify efficiently decodable codes that can attain or even exceed the hashing bound. A class of such codes will not only provide new approaches for obtaining tighter quantum channel capacity (QCC) lower bounds but also inspire proposals for more efficient strategies of error correction [qccapacity_very_noisy, degenerate_qu_code_pauli_channels, lower_bounds_capacity_birgitta, leditzky_pauli_hashing, xzzx_surface_code].
Using tensor-network decoding techniques from [farrelly_tn_codes, parallel_decoding_tn_codes], we study zero-rate versions of six different classes of holographic codes which are constructed from the seed codes of the
  1. (a)
    $\llbracket 5,1,3\rrbracket$, i.e. Harlow–Preskill–Pastawski–Yoshida (HaPPY) code [happy_paper]
  2. (b)
    Tailored $\llbracket 7,1,3\rrbracket$ code [tailored_713]
  3. (c)
    $\llbracket 7,1,3\rrbracket$ Steane code [harris_css]
  4. (d)
    $\llbracket 6,1,3\rrbracket$ code [farrelly_tn_codes]
  5. (e)
    Surface-code fragment (SCF) [harris_iod], and
  6. (f)
    A Clifford-deformed variant of the Steane code.
@@
Our main results can be stated as follows. In the limit of pure 2-Pauli noise channels, the thresholds calculated for zero-rate SCF and $\llbracket 6,1,3\rrbracket$ holographic codes exceed current state-of-the-art results [leditzky_pauli_hashing, lower_bounds_capacity_birgitta, qccapacity_very_noisy]. To our knowledge, no other class of quantum code has thus far been able to supersede the hashing limit for pure 2-Pauli noise. In fact, all of the holographic codes tested in the limit of pure 2-Pauli noise either come within $2\%$ of, attain, or overcome some portion of the hashing bound. Furthermore, all codes tested were found to either attain or overtake the hashing bound for some bias configuration of Pauli noise channel. Finally, we also apply the method of Clifford deformations from [3d_surface_hashing, clifford_deformed_surface] to the holographic Steane code, demonstrating that the threshold of the code can be drastically modified for depolarizing, pure X, and pure Z noise channels. Our results provide strong evidence that holographic codes possess high resilience against various types of Pauli noise channels and can be tuned accordingly. In sharp contrast with other code families, maximum likelihood decoding of holographic codes and the computation of their weight enumerator polynomials are possible in polynomial time, even with exact tensor network contractions [farrelly_tn_codes, cao2024quantum], further bolstering their practical potential for both quantum error correction and for numerical studies of QCC bounds.
@@
The remainder of this article is organized as follows: we review in the background (Section II) the basics of holographic codes and their properties for practical QEC in Section II.1, as well as an explanation for how tensor-network decoding functions (Section II.3). In Section III, we begin by discussing first the experimental setup for our simulations, the codes tested, and the biased-noise points considered (Section III.1); in Section III.2, we present all of the results of our study. Finally, we close with a discussion of future directions and implications of our work (Section IV), followed by a discussion of implications for quantum coding theory (Section IV.1), as well as practical directions for holographic codes (Section IV.2).
@@ II.1 Holographic Quantum Error Correction
Originally developed to model the AdS/CFT correspondence [witten, maldacena, klebanov, nastase, takayanagi_book], holographic quantum error correction involves encoding maps from bulk (logical) degrees of freedom on a d+1d{+}1-dimensional hyperbolic space to $d$-dimensional boundary (physical) degrees of freedom. While in full AdS/CFT the bulk and boundary degrees of freedom are associated with weakly-coupled quantum gravity on an asymptotically anti-de Sitter (AdS) background and strongly-coupled conformal field theory (CFT), respectively, aspects of this duality for $d=1$ can be captured by simple tensor network codes [almheiri2015bulk, happy_paper, Harlow:2018fse, holographic_codes_topical_review].
Figure 1: The holographic tensor-network codes considered in this work: (a) The Harlow-Preskill-Pastawski-Yoshida (HaPPY) code [happy_paper] whose pentagon-hexagon geometry also underlies the holographic surface-code fragment (SCF) model [harris_iod], (b) the holographic $\llbracket 6,1,3\rrbracket$ code [farrelly_tn_codes], and (c) the holographic Steane code [harris_css]. All are defined as tensor network contractions of copies of a fixed $q$-leg tensor on a hyperbolic tiling, with the central tensor used as a seed tensor for the encoding isometry of a $\llbracket q-1,1,d\rrbracket$ code with some distance $d$. The remaining tensors have all $q$ legs contracted in the plane, leading to a larger encoding map between one logical qubit (central red leg) and the boundary physical qubits (open black legs). Here we depict the codes with $R=2$ layers of edge inflation. In the $R\to\infty$ limit, the rate of each code goes to zero.
Holographic quantum codes are, at their core, generalized concatenated codes [poulin2005subsystem, lidar_qec, williamson_ft_logical_gates, gottesman_thesis_stabilizer_codes, heterogeneous_holo_qrm, baspin2024wire, liu2024subsystem], as their tree-like concatenation structure is a consequence of the bulk logical qubits being divisible into radial layers [happy_paper, conformal_quasi_holo, central_charges_jahn_eisert]. The most widely studied instances of holographic codes also fall into the category of stabilizer codes [happy_paper, harris_css, steinberg_evenbly_codes1], despite some restrictions on their representation of holographic dualities [Dong:2018seb, Akers:2018fow, Cao:2023mzo] and the possibility to construct more sophisticated non-stabilizer or even approximate holographic codes [hmera_approx, pollack2022understanding, cao2021approximate, kim2017entanglement, xp_stabilizer]. We shall therefore limit ourselves to stabilizer versions of holographic codes in the present work.
Stabilizer codes themselves are defined as codes for which the logical operators $\mathsf{L}$ and codeword stabilizers $\mathsf{S}$ consist of elements $P_{i}\in\mathsf{P}^{n}$, where $\mathsf{P}^{n}$ is the $n$-qubit Pauli group, $P_{i}$ takes the form $P_{i_{1}}\otimes\cdots\otimes P_{i_{n}}$, and $P\in\{I,X,Y,Z\}$, i.e., the single-qubit Pauli operators. It is also wholly possible to consider finite-rate versions of holographic codes, wherein individual seed tensors can be specified with logical-qubit implantation. In this work, we consider only zero-rate versions of holographic codes, wherein only the central logical index remains (Figure 1)  [^1] One can additionally construct subsystem versions of holographic quantum codes. Here, one chooses to subdivide the physical Hilbert space as $\mathsf{H}_{p}=\mathsf{H}\otimes\mathsf{\bar{H}}$, where $\mathsf{H}$ represents the code subspace. Then, given $\mathsf{H}$, we can decompose the code subspace further as $\mathsf{H}=\mathsf{H}_{L}\otimes\mathsf{H}_{G}$, where $\mathsf{H}_{L}$ and $\mathsf{H}_{G}$ represent the logical and gauge subsystems of the code subspace, respectively. Holographic codes have been previously treated as subsystem stabilizer codes in [cao2021approximate, williamson_ft_logical_gates, uberholography]. .
Represented as tensor networks, the bulk geometry of holographic codes is that of a regular hyperbolic tiling, wherein each $n$-gon tile is filled out by one tensor of $n$ planar (i.e., non-logical) legs. Starting from the central seed tensor whose single logical leg describes degrees of freedom in a small bulk patch in holography, the tensor network (and the bulk region it represents) can be iteratively extended via inflation steps, Here, each of these extensions can be realized by adding a layer of tensors (contracted) to the tensor network [conformal_quasi_holo, central_charges_jahn_eisert]. While this process is not unique, we will only apply edge inflation in this work (i.e. tensors/tiles are added to all open legs/edges of the previous step). We denote the number of inflation steps as $R$, with $R=0$ (no step) describing the seed tensor by itself, with Figure 1 showing holographic codes on various hyperbolic tilings at $R=2$.
@@
From a practical perspective, there are many reasons to find the concept of a holographic quantum code appealing. Firstly, holographic codes are relatively simple to construct and scale from small, atomized examples to larger example codes. As a direct consequence of the quantum LEGO formalism [cao_quantum_lego], concepts such as code structure and transversal logical operations are straightforward to intuit [heterogeneous_holo_qrm, cao2025growing]. Secondly, the boundary of holographic codes exhibits a quasiperiodic self-similarity [conformal_quasi_holo, central_charges_jahn_eisert], allowing one to rescale the number of physical and logical qubits using local inflation rules that also determine the code’s rate and distance scaling [parallel_decoding_tn_codes, evenbly_codes2], both of which fare better than other constructions [campbell2017roadstoward, terhal2015quantum, fowler2012surface, breuckmann_qldpc]. Thirdly, it is known that most seed tensors for holographic codes can be easily mapped to graph states, implying that efficient preparation schemes likely exist [engineering_holography_graph, raissi_graph1, raissi_graph2, helwig_ame_graph]. Lastly, and perhaps most consequentially, it has been shown that holographic codes exhibit high resilience against various noise channels. Indeed, it was first shown in [happy_paper] that holographic quantum error correction codes exhibit high thresholds against the quantum erasure channel; subsequently, several works have shown the capability of certain codes constructions to protect against depolarizing and 1-Pauli noise, potentially as well as topological codes [parallel_decoding_tn_codes, harris_css, harris_iod, farrelly_tn_codes, deconfinement, evenbly_codes2]. To our knowledge, a systematic biased-noise threshold study for even zero-rate holographic codes has not been performed, let alone for their constant-rate versions 111[cao2024quantum] discusses biased noise in finite-rate HaPPY pentagon codes, but it only captures the non-detectable error probability for a code of fixed length. . As such, our work takes the first step towards understanding holographic codes under more generalized noise channels.
@@
The main ingredient of a holographic quantum code is the seed tensor, defining the encoding map for a single tile (in Section II.2 we will provide descriptions of several seed codes which may be unfamiliar to the reader, such as the SCF and the tailored $\llbracket 7,1,3\rrbracket$ codes). Here we consider the zero-rate case with a logical qubit only on the central tile; on the remaining tiles all tensor legs are planar, effectively dispersing the logical information towards the tiling boundary. Note that in generic holographic codes, each tile may hold a logical qubit, leading to a nonzero asymptotic rate as the number $R$ of layers is increased.
Currently, several methods exist by which one may create a holographic code. Firstly, one may employ methods of gauge fixing for various logical qubits in the bulk. This step leaves each encoding tensor with only physical legs that act isometrically from one layer to the next. In the second case, one can tile the bulk by surrounding a code with parameters $\llbracket n,k,d\rrbracket$ by codes with parameters $\llbracket n+1,0,d+1\rrbracket$. In both cases, the contracted tensor network then forms one large encoding isometry from a single logical qubit to the physical boundary qubits given suitable (e.g. perfect [happy_paper]) tensors. Here, we opt for the second case. Thirdly, copies of the same seed tensor are typically used to construct such codes, but one may also employ several different seed tensors in similar spirit to heterogeneous concatenation methods for tree-style concatenated codes [concat1, concat2, heterogeneous_holo_qrm]. We note that all of these techniques it was recently shown that holographic codes are an instance of generalized code concatenation [generalized_code_concat, heterogeneous_holo_qrm].
II.2 Seed Tensors Used in this Paper
Here we provide information on the seed codes utilized in this paper: the perfect $\llbracket 5,1,3\rrbracket$ code [laflamme1996perfect]; the Steane code [Steane_1996]; the $\llbracket 6,1,3\rrbracket$ code [shaw2008encoding]; the surface-code fragment (SCF) from [harris_iod] and the tailored $\llbracket 7,1,3\rrbracket$ code [tailored_713].
Firstly, The famous perfect code of [laflamme1996perfect] is a $\llbracket 5,1,3\rrbracket$ code with cyclically permuted stabilizer support, as shown in Table 1. The perfect code is an example of generalized XZZX toric codes [tailored_513_xzzx_codes], and is known to exhibit high resilience against biased 1-Pauli noise [xzzx_surface_code, tailored_713]. Equally famous in its own right, the Steane code is the prototypical example of a Calderbank-Shor-Steane (CSS) code, in which two classical codes are utilized in order to construct a quantum code. The Steane code in particular is a self-dual CSS code, meaning that the X-/Z- portions of the stabilizer generators are exactly coinciding. More information on both codes can be found in [lidar_qec, nielsen_chuang], and the stabilizers and logical operators for the code can be found in Table 1.
The surface-code fragment (SCF) is the smallest surface code, and is known to possess a block-perfect encoding tensor, following the work of [harris_iod]. The stabilizers and logical operators for the SCF seed tensor are shown in Table 1. The block-perfect condition is taken to mean that, given any permutation of cyclic permutations for the SCF tensor indices $\{1,2,3,4,\bar{6},5\}$ (where $\bar{6}$ denotes the logical index), we can construct the condition whereby
[…]
where $|A|=\lfloor\frac{n}{2}\rfloor$. The SCF seed tensors are an instance of planar maximally-entangled (PME) states [doroudiani2020planar], meaning that only adjacent groupings into bipartitions can reduce to maximally mixed states.
The tailored $\llbracket 7,1,3\rrbracket$ code is a generalization of the celebrated perfect code [laflamme1996perfect] and was first introduced in [tailored_713]. The stabilizers and logical operators for the code are listed in Table 1. As stated in [tailored_513_xzzx_codes, wen2003quantum], both the perfect code and the tailored $\llbracket 7,1,3\rrbracket$ codes can be regarded as the smallest toric codes with so-called shifted boundary conditions. The tailored $\llbracket 7,1,3\rrbracket$ code was shown in [tailored_713] to exhibit better threshold performance than the Steane code in the setting of highly-biased noise.
II.3 Tensor-Network Decoding